Consider the following equilibrium: PbI2(s)↽−−⇀Pb2+(aq)+2I−(aq) If the concentration of the Pb2+ ion at a certain point is 2.1×10−3 M, and Ksp=7.5×10−11, will a precipitate form? Select the correct answer below: yes, because Q>Ksp yes, because QKsp no, because Q
Answer #1

[Pb2+] = 2.1*10^-3 M

PbI2 <-> Pb2+ + 2 I-

s 2s

So,

[I-] = 2s

= 2*2.1*10^-3 M

= 4.2*10^-3 M

Use:

Qsp = [Pb2+][I-]^2

= (2.1*10^-3) * (4.2*10^-3)^2

= 8.8*10^-6

Qsp > ksp, So, precipitate will form

Answer:

yes, because Q>Ksp